OpenStudy (ray10):

Find the first three nonzero terms of the Maclaurin expansion of the function; \[\large f(x)=8*\cos(x)\]

OpenStudy (mimi_x3):

have you tried it?

OpenStudy (ray10):

sure have and I seem to fail almost immediately :( I just don't understand where to go with it

OpenStudy (mimi_x3):

so, x = 0 f(x)= 8cos(x) => f(0) = 8cos(0) = 8 f'(x) = -8sin(x) => f'(0) = -8(0 = 0 f''(x) = -8cos(x) => f''(0) = -8(1) = -8

OpenStudy (mimi_x3):

have you done something like that?

OpenStudy (mimi_x3):

oh shoot, is macaluarin expansion and series the same lol?

OpenStudy (ray10):

Ah Maclaurin is a different form of test is what I am aware of. My notes say that where a function and all its derivatives are computed where x=0, we then are dealing with a Maclaurin series :S

OpenStudy (mimi_x3):

ok, so im on the right track :D

OpenStudy (ray10):

yaaaay :D good to hear

OpenStudy (mimi_x3):

continuing with the above f'''(x) = 8sin(x) => f'''(0) = 0 \[ f(x) = f(x) + \frac{f'(x)}{1!}x + \frac{f''(x)}{2!}x^2 + \frac{f'''(x)}{3!}x^3+...\]

OpenStudy (ray10):

so being nonzero terms we must use each derivative where x=0?

OpenStudy (mimi_x3):

so subbing it in \[ f(x) = 8 + \frac{0}{1!}x + \frac{-8}{2!}x^2 +\frac{0}{3!}x^3 + \frac{8}{4!}x^4+..\] \[ f(x) = 8 + -\frac{8}{2!}x^2+ \frac{8}{4!}x^4+..\]

OpenStudy (mimi_x3):

non-zero simply means to not include the terms that have a \(0\), if that make sense.

OpenStudy (mimi_x3):

Well, it would be better, if you have the answer for it. Don't always trust me.

OpenStudy (ray10):

I understand that now :). Yes but you helped me out in understanding it! I wish I had then answers for it now, at least the solutions to see. But because I didn't know it I thought I would ask :)

OpenStudy (mimi_x3):

alrighty, good luck :) i confirmed it with wolf, and my solution is right, so dw :) http://www.wolframalpha.com/input/?i=maclaurin+exansion+8cos%28x%29

OpenStudy (ray10):

I had a look, but where does it show the answer for it? :)

OpenStudy (mimi_x3):

http://gyazo.com/70c331395db6c26ccd6e2d85be403d58

OpenStudy (mimi_x3):

exact same answer, of what i got, but i was too lazy to simplify it, lol

OpenStudy (ray10):

oh haha well thank you @Mimi_x3 !! :D

OpenStudy (mimi_x3):

No worries :)